OpenStudy (anonymous):

Suppose the derivative of f exists, and assume that f(5)=3, and f′(5)=5. Let g(x)=x^2f(x), and h(x)=f(x)/x−4. a) g′(5)=______________ Find the equation of the tangent line to g(x) at x=5. y=_____________ b) h′(5)=______________ Find the equation of the tangent line to h(x) at x=5. y=__________________

OpenStudy (anonymous):

I know h'(5)= 2 but i'm stumped on the remaining parts

OpenStudy (anonymous):

g'(x) = 2xf(x) + x^2f'(x) can you do it now?

OpenStudy (anonymous):

is g′(5) = 155 ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@Coolsector oh i did'nt see ur post

OpenStudy (anonymous):

that's the right way

OpenStudy (anonymous):

why is g'(x) = 2xf(x) + x^2f'(x)? I understand the first part but I dont know where the second part of the equation came in

OpenStudy (anonymous):

using the product rule : g(x) = x^2 * f(x) g'(x) = (x^2)' * f(x) + (x^2) * f'(x)

OpenStudy (anonymous):

in general : (uv)' = u'v + v'u

OpenStudy (anonymous):

oh wow I dont know why I didnt see that. Thank you!! :)

OpenStudy (anonymous):

yw